# Study of Flat Spectrum Radio Quasars and BL Lacertae Objects as Sources of Diffusive Ultra High-Energy Cosmic Rays

### v1

@@ V Summary and Conclusion
The full-sky flux distributions constructed using the HEALPix formalism reveal distinct anisotropy patterns across different energy scales and source classes. In the FSRQ case, anisotropy becomes increasingly pronounced at higher energies ($50$ EeV), highlighting the dominance of nearby high-luminosity sources and reduced contribution from distant ones due to energy loss effects. Conversely, BL Lac sources display a relatively smoother and more isotropic flux pattern, consistent with their lower luminosity and uniform spatial distribution. These results emphasize the combined influence of source properties, spatial evolution, and propagation effects in shaping the observed UHECRs flux.
The positional cross-matching of identified hotspot regions with the SIMBAD database further supports a possible association between localized flux enhancements and nearby extragalactic sources. In particular, the hotspot near $(\mathrm{RA}=163.42^{\circ},,\mathrm{Dec}=-88.26^{\circ})$ coincides with several AGNs and galaxies as given in Table 2 and 3. Such correlations indicate that these extragalactic sources could contribute to the observed high-energy flux excess.
Overall, the present analysis demonstrates that the mixed-composition models under this method can reproduce key spectral and compositional features observed in UHECRs data. The anisotropy signatures and hotspot correlations offer potential clues to the origin of UHECRs, motivating further multi-messenger and directional studies with improved statistics and higher-resolution datasets.

### v2

@@ V Summary and Conclusion
We constructed full-sky maps of the integral UHECR flux in the energy range $10\leq E\leq 100~\mathrm{EeV}$ using the HEALPix framework with an angular resolution of ${\tt Nside}=256$. The maps were generated assuming a uniform comoving distribution of synthetic sources for each source class, without using any observational source catalogue. Source positions were assigned randomly over the full sky, and each source contribution was added to the corresponding HEALPix pixel. The integral flux above the chosen energy threshold was obtained by numerically integrating the propagated spectrum up to the maximum injection energy. To suppress small-scale numerical fluctuations and highlight large-scale anisotropy, the maps were smoothed with a Gaussian function with a width of $2^{\circ}$. These skymaps are intended for illustration only, as the underlying source distribution is entirely synthetic. At energies above $\sim 10~\mathrm{EeV}$, the limited propagation horizon implies that only a few nearby sources dominate the flux, naturally producing localized excesses or “hotspots”. Such features arise from proximity effects in the simulations and do not represent real astrophysical anisotropies.
To provide contextual information on the sky regions corresponding to the synthetic hotspots, we performed a positional cross-matching with the SIMBAD astronomical database using the astroquery.simbad module. This analysis was carried out solely as an exploratory and orientational exercise. We emphasise that the skymaps themselves contain no observational directional information, and the subsequent comparison with catalogued galaxies or active galactic nuclei does not imply any physical association or source identification. The cross-matching serves only to illustrate which known extragalactic objects happen to lie near the directions of the model-generated flux enhancements. The identified nearby galaxies and AGN-like sources are summarised in Tables 2 and 3.
Overall, our results demonstrate how energy-dependent magnetic propagation effects and finite CRs horizons can naturally produce anisotropy like features in UHECR skymaps, even for uniformly distributed source populations. The framework developed here provides a self–consistent basis for interpreting such features in synthetic models and for exploring how different source classes may contribute to the observed UHECR sky.